OpenStudy (zzr0ck3r):

Convex analysis Does anyone know anything about projected subgradient/gradient algorithm to solve linear programming problems?

OpenStudy (zzr0ck3r):

we need to use the gradient algorithm with step size a_k = 1/sqrt(k) to solve the following problem maximize 5x+4y x<=6 .25x+y<=6 3x+2y<=22 x,y>=0